# An efficient likelihood-free Bayesian inference method based on sequential neural posterior estimation

### v3

@@ Abstract
Sequential neural posterior estimation (SNPE) techniques have been recently proposed for dealing with simulation-based models with intractable likelihoods. Unlike approximate Bayesian computation, SNPE techniques learn the posterior from sequential simulation using neural network-based conditional density estimators by minimizing a specific loss function. The SNPE method proposed by Lueckmann et al. (2017) used a calibration kernel to boost the sample weights around the observed data, resulting in a concentrated loss function. However, the use of calibration kernels may increase the variances of both the empirical loss and its gradient, making the training inefficient. To improve the stability of SNPE, this paper proposes to use an adaptive calibration kernel and several variance reduction techniques. The proposed method greatly speeds up the process of training and provides a better approximation of the posterior than the original SNPE method and some existing competitors as confirmed by numerical experiments. We also manage to demonstrate the superiority of the proposed method for a high-dimensional model with real-world dataset.
@@ 2 Background
where $Z(x,\phi)=\int q_{F(x,\phi)}(\theta)\tilde{p}(\theta)/p(\theta)\mathrm{d}\theta$ represents a normalization constant. Following [42, Proposition 1], when $\tilde{q}_{F(x,\phi)}(\theta)$ converges to the proposal posterior $\tilde{p}(\theta|x)\propto\tilde{p}(\theta)p(x|\theta)$, it implies that $q_{F(x,\phi)}(\theta)$ converges to $p(\theta|x)$. To compute $Z(x,\phi)$, [42] proposes the atomic method, which enables the analytical computation of $Z(x,\phi)$ by setting $\tilde{p}(\theta)$ as a discrete distribution. However, this method necessitates nested simulation when computing $Z(x,\phi)$, which implies extra simulation and computational costs.
@@ 3.2.2 Sample recycling and multiple importance sampling
This is an application of [53, Theorem 1]. ∎
@@ 3.2.3 Adaptive calibration kernel based on an ESS criterion
In the absence of MISR, the number of samples used in each round remains constant. A common criterion for the selection of ESS in this scenario is to define it as a fraction of the current iteration sample size $N$, that is, $\mathrm{ESS}=\gamma N$ for $\gamma\in(0,1)$. In each round, after simulating the dataset $\{(\theta_{i},x_{i})\}_{i=1}^{N}$, the value of $\tau$ is numerically determined by solving the equation $\mathrm{ESS}=\gamma N$ using the bisection method. This approach achieves a balance between the bias and variance of the estimator.
@@ 4 Experiments
In this section, we conduct a series of comparative analyses of our proposed strategies with existing methods to demonstrate their effectiveness. Since our strategies are built upon the SNPE-B method, hence we begin with an ablation study. Subsequently, we then compare our methods with other likelihood-free Bayesian computation methods such as SNPE-A [42], APT [20] and SNL [44].
@@
4.1 Ablation study on the SNPE-B method
@@
Figure 2: Ablation experiments on proposed strategies. A. Performance on M/G/1 queuing model. B. Performance on Lotka-Volterra predation model. The horizontal axis represents the round of training and the error bars represent the mean with the upper and lower quarterlies. Our proposed method exhibits better performance compared to the original method.
In this section, we provide a detailed comparison of the performance of our proposed strategies, including adaptive calibration kernels (ACK), defensive sampling (DS), and multiple importance sampling and recycling (MISR) methods with the standard SNPE-B method. In some cases, the support of the ground truth posterior is bounded, while the neural posterior always has an unbounded support. To circumvent this, we consider two strategies to account for the inconsistency of supports. The first one is known as truncated posterior (TP), in which we directly truncate the neural posterior after training, this method suffers from low sampling efficiency and leakage of mass. The second one is the parameter space transformation (PST), which is detailed in Section 3.3. To study the difference between TP and PST, we additionally examine these two cases with all other strategies in our experiment, which is represented as ‘All-SNPE-B (TP)’ and ‘All-SNPE-B (PST)’, respectively. ‘SPNE-B’ in the following figures stands for the standard SNPE-B method without taking any strategies we proposed.
Figure 2 shows the results of the ablation experiments. The utilization of the ACK method effectively reduces LMD, which is predictable since the calibration kernel directly assigns weight to the sample pair $(\theta,x)$ based on the distance between $x_{o}$ and $x$. Nevertheless, it does not perform well on other metrics and still suffers from unstable learning. Surprisingly, the introduction of the calibration kernel can sometimes exacerbate this instability. DS and MISR indeed reduce the variance and thus narrow the error bar. However, they tend to lead to a loss in performance. There are two reasons for this occurrence. On the one hand, defensive distribution brings in irrelevant samples, making the training samples not sufficiently concentrated. On the other hand, MISR is prone to including ‘bad’ samples from earlier rounds, which also results in lower sample quality.
@@
4.2 Compare to other likelihood-free inference methods
@@
We refer the reader to Figure 1 for a clear understanding of the advantages of All-SNPE-B in seeking modes. We assign this credit to ACK since it helps the neural network to recognize the importance of samples according to the distance between $\mathbf{x}_{o}$ and its observation. We report a better approximation around the mode and the ground-truth distribution with the marginal plot. A similar result is also observed in the case of ACK-APT, which shows consistency instead of divergence. For example, Figure 1-B demonstrates the significant improvement of ACK over APT.
@@
Finally, we validate our findings with a simulated data. We fix the true model parameters and generate observations $x_{o}$ accordingly, enabling the calculation of the NLOG metric. The simulation period was set to 4 weeks, with $T=28$, resulting in a summary statistic dimensionality of $d=112$. In Figure 5-B, we present the NLOG and LMD values for each method. Additionally, we report the posterior marginal density plots over 10 runs in Figure 6-B. Our method successfully captures the true model parameters, whereas other methods fail (notably, including APT) in this scenario. The results show that the All-SNPE-B method achieves consistently stable and remarkable performance, while the ACK-APT method, though less stable, still outperforms the other baseline methods. This demonstrates the power of the ACK strategy, which is designed to help the model to identify effectively which training samples within the high-dimensional parameter space deserve greater focus, rather than attempting to approximate the entire posterior distribution across the full space.
@@ 5 Discussion
It is worth mentioning that the proposed form of calibration kernels in SNPE-B [35], is adaptive since it depends on the Gaussian proposal distribution, which updates with rounds. Our proposed method is also adaptive since tne calibration kernel rate updates according to solving an equation related to ESS.
@@ Appendix C Additional model information
This section provides detailed information on the SLCP and Gaussian linear models.
@@
$E_{n}=(1,0,\cdots,0)$ is a $n$-dimensional vector, and $\delta(t)$ is an indicator function epresenting the summer season. The evolution matrix $G_{t}^{(i)}$ comprises several components: a block $J_{2}$ to model the local linear trend of the latent level, a permutation matrix $P_{6}$ to capture the weekly seasonal effect, and specific terms to describe the summer effect using $\theta_{9,t}^{(i)}$. The complete model is constructed by combining $F_{t}^{(i)}$ and $G_{t}^{(i)}$ with the Kronecker product, resulting in $F_{t}=F_{t}^{(i)}\otimes I_{4}$ and $G_{t}=G_{t}^{(i)}\otimes I_{4}$, with $\theta_{t}=(\theta_{t}^{(1)},\ldots,\theta_{t}^{(4)})$. The process noise is defined as $w_{t}\sim N(0,10^{-5}I_{36})$, and the initial state is set as $\theta_{0}\sim N(0,10^{-1}I_{36})$.
@@
Figure 7: Ablation experiments on proposed strategies. A. Performance on the SLCP model. B. Performance on the Gaussian linear model. The horizontal axis represents the round of training and the error bars represent the mean with the upper and lower quartiles.

### v4

@@ Abstract
Sequential neural posterior estimation (SNPE) techniques have been recently proposed for dealing with simulation-based models with intractable likelihoods. Unlike approximate Bayesian computation, SNPE techniques learn the posterior from sequential simulation using neural network-based conditional density estimators by minimizing a specific loss function. The SNPE method proposed by Lueckmann et al. (2017) used a calibration kernel to boost the sample weights around the observed data, resulting in a concentrated loss function. However, the use of calibration kernels may increase the variances of both the empirical loss and its gradient, making the training inefficient. To improve the stability of SNPE, this paper proposes to use an adaptive calibration kernel and several variance reduction techniques. The proposed method greatly speeds up the process of training and provides a better approximation of the posterior than the original SNPE method and some existing competitors as confirmed by numerical experiments. We also managed to demonstrate the superiority of the proposed method for a high-dimensional model with a real-world dataset.
@@ 2 Background
where $Z(x,\phi)=\int q_{F(x,\phi)}(\theta)\tilde{p}(\theta)/p(\theta)\mathrm{d}\theta$ represents a normalization constant. Following [43, Proposition 1], when $\tilde{q}_{F(x,\phi)}(\theta)$ converges to the proposal posterior $\tilde{p}(\theta|x)\propto\tilde{p}(\theta)p(x|\theta)$, it implies that $q_{F(x,\phi)}(\theta)$ converges to $p(\theta|x)$. To compute $Z(x,\phi)$, [43] proposes the atomic method, which enables the analytical computation of $Z(x,\phi)$ by setting $\tilde{p}(\theta)$ as a discrete distribution. However, this method necessitates nested simulation when computing $Z(x,\phi)$, which implies extra simulation and computational costs.
@@ 3.2.2 Sample recycling and multiple importance sampling
This is an application of [54, Theorem 1]. ∎
@@ 3.2.3 Adaptive calibration kernel based on an ESS criterion
In the absence of MISR, the number of samples used in each round remains constant. A common criterion for the selection of ESS in this scenario is to define it as a fraction of the current iteration sample size $N$, that is, $\mathrm{ESS}=\gamma N$ for $\gamma\in(0,1)$. In each round, after simulating the dataset $\{(\theta_{i},x_{i})\}_{i=1}^{N}$, the value of $\tau$ is numerically determined by solving the equation $\mathrm{ESS}=\gamma N$ using the bisection method [12]. This approach achieves a balance between the bias and variance of the estimator.
@@ 4 Experiments
In this section, we conduct a series of comparative analyses of our proposed strategies with existing methods to demonstrate their effectiveness. Since our strategies are built upon the SNPE-B method, hence we begin with a simulation study. Subsequently, we then compare our methods with other likelihood-free Bayesian computation methods such as SNPE-A [43], APT [21] and SNL [45].
@@
4.1 Simulation study on the SNPE-B method
@@
Figure 2: Simulation experiments on proposed strategies. A. Performance on M/G/1 queuing model. B. Performance on Lotka-Volterra predation model. The horizontal axis represents the round of training and the error bars represent the mean with the upper and lower quarterlies. Our proposed method exhibits better performance compared to the original method.
In this section, we provide a detailed comparison of the performance of our proposed strategies, including adaptive calibration kernels (ACK), defensive sampling (DS), and multiple importance sampling and recycling (MISR) methods with the standard SNPE-B method. In some cases, the support of the ground truth posterior is bounded, while the neural posterior always has an unbounded support. To circumvent this, we consider two strategies to account for the inconsistency of support. The first one is known as truncated posterior (TP), in which we directly truncate the neural posterior after training, this method suffers from low sampling efficiency and leakage of mass. The second one is the parameter space transformation (PST), which is detailed in Section 3.3. To study the difference between TP and PST, we additionally examine these two cases with all other strategies in our experiment, which are represented as ‘All-SNPE-B (TP)’ and ‘All-SNPE-B (PST)’, respectively. ‘SPNE-B’ in the following figures stands for the standard SNPE-B method without taking any strategies we proposed.
Figure 2 shows the results of the simulation experiments. The utilization of the ACK method effectively reduces LMD, which is predictable since the calibration kernel directly assigns weight to the sample pair $(\theta,x)$ based on the distance between $x_{o}$ and $x$. Nevertheless, it does not perform well on other metrics and still suffers from unstable learning. Surprisingly, the introduction of the calibration kernel can sometimes exacerbate this instability. DS and MISR indeed reduce the variance and thus narrow the error bar. However, they tend to lead to a loss in performance. There are two reasons for this occurrence. On the one hand, defensive distribution brings in irrelevant samples, making the training samples not sufficiently concentrated. On the other hand, MISR is prone to including ‘bad’ samples from earlier rounds, which also results in lower sample quality.
@@
4.2 Comparisons with other likelihood-free inference methods
@@
We refer the reader to Figure 9 for a clear understanding of the advantages of All-SNPE-B in seeking modes. We assign this credit to ACK since it helps the neural network recognize the importance of samples according to the distance between $\mathbf{x}_{o}$ and its observation. We report a better approximation around the mode and the ground-truth distribution with the marginal plot. A similar result is also observed in the case of ACK-APT, which shows consistency instead of divergence. For example, Figure 9-B demonstrates the significant improvement of ACK over APT.
@@
Finally, we validate our findings with simulated data. We fix the true model parameters and generate observations $x_{o}$ accordingly, enabling the calculation of the NLOG metric. The simulation period was set to 4 weeks, with $T=28$, resulting in a summary statistic dimensionality of $d=112$. In Figure 5-B, we present the NLOG and LMD values for each method. Additionally, we report the posterior marginal density plots over 10 runs in Figure 6-B. Our method successfully captures the true model parameters, whereas other methods fail (notably, including APT) in this scenario. The results show that the All-SNPE-B method achieves consistently remarkable performance, while the ACK-APT method, though less stable, still outperforms the other baseline methods. The challenges faced by SNPE-B and APT arise from the high dimensionality of the sample space, even after incorporating summary statistics. In such settings, conditional density estimators struggle to focus on the sparse and highly informative training samples near $x_{o}$. The ACK strategy addresses this by prioritizing these critical regions, rather than attempting to approximate the entire posterior distribution across the full parameter space.
@@ 5 Discussion
It is worth mentioning that the proposed form of calibration kernels in SNPE-B [36], is adaptive since it depends on the Gaussian proposal distribution, which updates with rounds. Our proposed method is also adaptive since the calibration kernel rate updates according to solving an equation related to ESS.
@@ Appendix C Additional model information
This section provides detailed information on the SLCP model, Gaussian linear model, and g-and-k based state-space model.
@@
$E_{n}=(1,0,\cdots,0)$ is a $n$-dimensional vector, and $\delta(t)$ is an indicator function epresenting the summer season. The evolution matrix $G_{t}^{(i)}$ comprises several components: a block $J_{2}$ to model the local linear trend of the latent level, a permutation matrix $P_{6}$ to capture the weekly seasonal effect, and specific terms to describe the summer effect using $\theta_{9,t}^{(i)}$. The complete model is constructed by combining $F_{t}^{(i)}$ and $G_{t}^{(i)}$ with the Kronecker product, resulting in $F_{t}=F_{t}^{(i)}\otimes I_{4}$ and $G_{t}=G_{t}^{(i)}\otimes I_{4}$, with $\theta_{t}=(\theta_{t}^{(1)},\ldots,\theta_{t}^{(4)})$. The process noise is defined as $w_{t}\sim N(0,10^{-5}I_{36})$, and the prior of the initial state is set as $\theta_{0}\sim N(0,10^{-1}I_{36})$.
@@
Figure 7: Simulation experiments on proposed strategies. A. Performance on the SLCP model. B. Performance on the Gaussian linear model. The horizontal axis represents the round of training and the error bars represent the mean with the upper and lower quartiles.